# On Partition Functions for Time-Inhomogeneous Branching Random Walks

### 15 Mar 2026 v1

@@ 3 Proof of Main Results
Before we dive into the proof, we would like to illustrate a standard summation argument (the $L^{2}$-computation) used in our analysis, which usually forms the main technicality in various research in branching processes, and will be used repeatedly (in various forms) throughout the paper. For simplicity, we would first focus on our discussion in the time-homogeneous weights, and it would be straightforward to extend the computation to the time-inhomogeneous weights. The key idea is to fix a branch $\mathbf{|x|}=n$ and sum over all possible depths for $|\mathbf{y}|=n$ at which the pair of branches $\mathbf{x}$ and $\mathbf{y}$ diverge.
Within the BRWs framework, the covariance $\text{Cov}\left(H_{n}(\mathbf{x}),H_{n}(\mathbf{y})\right)$ for two branches $|\mathbf{x}|=|\mathbf{y}|=n$ depends exclusively on their branching depth $h\in\{0,\dots,n\}$—the depth at which they first diverge. Formally, for branches $\mathbf{x}=(x_{1},\dots,x_{n})$ and $\mathbf{y}=(y_{1},\dots,y_{n})$, define their coincidence depth as:
@@
with the convention that $C(\mathbf{x},\mathbf{y})=0$ if the set is empty. For such pair of branches, the covariance of their Hamiltonian satisfies $\text{Cov}\left(H(\mathbf{x}),H(\mathbf{y})\right)=2h$.
We fix a branch $\mathbf{|x|}=n$ (noting there are expected $d^{n}$ such branches) and sum over all possible branching depths $h\in\{0,\dots,n\}$ relative to $\mathbf{x}$. For each fixed $\mathbf{x}$ and depth $h$, the number of branches $\mathbf{y}$ satisfying $C(\mathbf{x},\mathbf{y})=h$ has expected value $d^{n-h}$, corresponding to the number of branches in a sub-Galton-Watson tree with (nh)(n-h) generations.
@@
Since the Laplace transform of $H_{n}(\mathbf{x})+H_{n}(\mathbf{y})$ for the sum of Hamiltonian for two branches $\mathbf{x},\mathbf{y}$ where $C(\mathbf{x},\mathbf{y})=h$ is given by:
$\mathbb{E}\left[e^{\beta(H_{n}(\mathbf{x})+H_{n}(\mathbf{y}))}\right]=\mathbb{E}\left[e^{2\beta(\tilde{H}(\mathbf{x}))}\right]\mathbb{E}\left[e^{\beta(H^{\prime}(\mathbf{x}))}\right]\mathbb{E}\left[e^{\beta(H^{\prime}(\mathbf{y}))}\right]$
where $\tilde{H}(\mathbf{x})$ has law $\mathcal{N}(0,h)$ and $H^{\prime}(\mathbf{x}),H^{\prime}(\mathbf{y})$ both have law $\mathcal{N}(0,(n-h))$ and they are all independent, so this equals to:
$\mathbb{E}\left[e^{\beta(H(\mathbf{x})+H(\mathbf{y}))}\right]=e^{2\beta^{2}h}(e^{\frac{\beta^{2}}{2}(n-h)})^{2}=e^{\beta^{2}(n+h)}$
As a result, the previous $L^{2}$ norm of $W_{n}^{\beta}$ equals to:
$\mathbb{E}\left[(W_{n}^{\beta})^{2}\right]=\sum_{h=0}^{n}d^{-h}e^{-n\beta^{2}}e^{\beta^{2}(n+h)}=\sum_{h=0}^{n}e^{h(\beta^{2}-\log d)}$
@@
(Time-inhomogeneous $L^{2}$ Critical Parameter $\bar{\beta}_{2}$)
@@
where $C$ is a constant depending only on $\beta$. For any fixed $\beta>\sqrt{\log d}$, we can choose $\delta>0$ sufficiently small such that $\log d+\beta^{2}(1-\delta)>0$. By the continuity of $f$, $\delta$ can be made arbitrarily small by taking $\epsilon$ small enough, and the series diverges as $n\to\infty$. ∎
@@
Theorem 3.3.
@@
We prove the $L^{1}$-convergence by dividing the branches into ”good environment” and its complement, for both time-homogeneous and time-inhomogeneous environments. Using the Cameron-Martin-Girsanov formula mentioned previously, we can see that the branches in ”good environment” in both cases dominates the total mass (i.e. the $L^{1}$-norm), and converges in $L^{2}$ to the same random variable.
Let $\beta<\alpha<2\beta$ be a fixed constant. For any branch $|\mathbf{x}|=n$, we define the ”good environment” event for the time-homogeneous BRWs, denoted $G_{n}^{\alpha}(\mathbf{x})$ (roughly speaking, it corresponds to the environment where the Hamiltonian of a branch $\mathbf{x}$ remains ”controlled”) for all $T\in\{n_{0},\ldots,n\}$ ,with $n_{0}\geq 0$ a fixed integer, by:
@@
which tends to 0 as $n_{0}\to\infty$ uniformly in $n$. A similar argument yields $\mathbb{E}[\bar{J}_{n}]=1-\epsilon^{\prime}(n_{0})$ where $\epsilon^{\prime}(n_{0})\to 0$ as $n_{0}\to\infty$ uniformly in $n$. While the right-hand side of (3) converges for any $\alpha>0$, we later use the good environment $G_{n}^{\alpha-\beta}(\mathbf{x})$ with $\alpha<2\beta$, justifying our restriction to $\beta<\alpha<2\beta$.
Next, we will show that when $\beta<\sqrt{2}\bar{\beta_{2}}$, $J_{n}$ and $\bar{J}_{n}$ are both uniformly bounded in $L^{2}$. Since $\bar{\beta}_{2}=\beta_{2}$, we have that $\bar{J}_{n}$ is uniformly bounded in $L^{2}$ when $\beta<\beta_{c}$. For pairs $(\mathbf{x},\mathbf{y})$ with $C(\mathbf{x},\mathbf{y})=h$, define the transformed measure $\tilde{\mathbb{P}}^{h}$ by:
@@
letting $\alpha\to\beta$, we can see that the series is uniformly bounded in $n$ when $\beta<\sqrt{2}\bar{\beta_{2}}$, by a direct comparison with the $L^{2}$ computation of $\bar{W}_{n}^{\beta}$ in equation 2. A similar comparison with the $L^{2}$ computation of $W_{n}^{\beta}$ yields that when $\beta<\sqrt{2}\beta_{2}=\beta_{c}$, $J_{n}$ is uniformly bounded in $L^{2}$.
@@
Fix $n_{1}>0$; we decompose $\mathbb{E}\left[(\bar{J}_{n}^{\beta})^{2}-\bar{J}_{n}^{\beta}J_{n}^{\beta}\right]$ by $C(\mathbf{x},\mathbf{y})=h\in\{0,...,n\}$ into two sums over $h\leq n_{1}$ and $h>n_{1}$, so that $\mathbb{E}\left[(\bar{J}_{n}^{\beta})^{2}-\bar{J}_{n}^{\beta}J_{n}^{\beta}\right]$ equals:
@@
For $\beta<\beta_{c}$ and any fixed $n_{1}>0$, form Lemma 3.4 we can see that $J_{n}$ and $\bar{J}_{n}$ converges to the same random variable $L^{2}$, since $\mathbb{E}[(\bar{J}_{n})^{2}]$ and $\mathbb{E}[\bar{J}_{n}J_{n}]$ has the following expansions:
@@
3.0.1 Remark:
It follows from the proof that such convergence holds jointly for finitely many functions $f$ that satisfy the properties specified in the present paper. However, this proof strategy does not automatically extend to the uniform convergence in such space of functions. As a counterexample, take the sequence of functions $(f_{n})_{n\geq 1}$ defined by $f_{n}(x)=(1-x)^{n}$ for all $n\geq 1$.
@@ 3.0.1 Remark:
Under the notation of Theorem 3.3’s proof, $J_{n}$ converges in $L^{2}$ for each fixed $n_{0}\geq 0$.
@@
Let $n_{1}>0$ be a fixed integer. For all $h\in\{0,\ldots,n_{1}\}$, the convergence
@@
The following proposition establishes the upper bound for $\bar{\beta}_{c}$, i.e. when $\beta>\beta_{c}$, $\bar{W}_{n}^{\beta}$ converges to 0 in probability. Together with Theorem 3.3, this confirms that the normalized partition function $\bar{W}_{n}^{\beta}$ exhibits distinct phase transitions in its limit in probability. Furthermore, the critical parameter $\bar{\beta}_{c}$ is defined by:
@@
exists and coincides with that for homogeneous weights, i.e., $\beta_{c}=\bar{\beta}_{c}$. This result underscores the insensitivity to environmental variations, indicating that the macroscopic behavior of the partition function is dominated by the ”early generations” of the branching structure, rather than being affected by prescribed variations in the branch weights.
Theorem 3.6.
@@
We define a transformed (random) probability measure $\bar{\mathbb{P}}^{n_{1}}$ by (where $\mathbf{P}$ is the law of the BRWs):
$\frac{\mathrm{d}\bar{\mathbb{P}}^{n_{1}}}{\mathrm{d}\mathbf{P}}=\frac{e^{-\frac{1}{2}\beta^{2}\sum_{i=1}^{n_{1}}f\left(\frac{i}{n}\right)+\beta\bar{H}_{n_{1}}(\mathbf{x})}}{\bar{W}_{n}^{n_{1}}}$ (11)
@@
For any function $F:D_{n_{1}}\to\mathbb{R}$ (recall that $D_{n_{1}}$ is the set of branches $|\mathbf{x}|=n$), we have
@@
which is $\mathcal{F}_{n_{1}}$-measurable (recall that $\mathcal{F}_{n_{1}}$ is the $\sigma$-field generated by the GW tree and the weights up to generation $n_{1}$).
It follows that
@@
By the Cauchy-Schwarz inequality, this implies
@@
To bound $\mathbb{E}\left[(\bar{W}_{n}^{n_{1}})^{2\alpha}\right]$, we will show it by computing its fractional moment:
$\mathbb{E}\left[(\bar{W}_{n}^{n_{1}})^{2\alpha}\right]\leq d^{n_{1}}\mathbb{E}\left[d^{-2\alpha n}e^{-\alpha\beta^{2}\sum_{i=1}^{n_{1}}f\left(\frac{i}{n}\right)+2\alpha\beta\bar{H}_{n_{1}}(\mathbf{x})}\right],$
which simplifies to:
@@
Using the strategy of the proof of these results, we end up this section by providing a novel proof for Theorem 2.1. Since the proof follows from the similar lines with that of Theorem 3.3, we prefer not to state all the details. Indeed, we can see that the proof is much simpler compared to the proof of Theorem 3.3: since $(W_{n}^{\beta})_{n\geq 0}$ defines a martingale, to show $L^{1}$ convergence it is suffice to show that it is uniformly integrable (i.e. $J_{n}$ is uniformly bounded in $L^{2}$ for each fixed $n_{0}>0$). The upper bound follows from the same argument as the last part of the proof of Proposition 3.6: for $\alpha\in(0,1)$ we have:
@@
3.1 Behavior of Partition Function at Criticality $\bar{\beta}_{c}$

### 27 Mar 2026 v2

@@ 3 Proof of Main Results
3.1.1 Lower Bound for Time-inhomogeneous Critical Parameter
The lower bound follows from a similar idea in the construction of sub-critical regime of GMC measure in [4]. Before we dive into the proof, we would like to illustrate a standard summation argument (the $L^{2}$-computation) used in our analysis, which usually forms the main technicality in various research in branching processes, and will be used repeatedly (in various forms) throughout the paper. For simplicity, we would first focus on our discussion in the time-homogeneous weights, and it would be straightforward to extend the computation to the time-inhomogeneous weights, by independence of increments. The key idea is to fix a branch $\mathbf{|x|}=n$ and sum over all possible depths for $|\mathbf{y}|=n$ at which the pair of branches $\mathbf{x}$ and $\mathbf{y}$ diverge.
Within the BRWs framework, the covariance $\text{Cov}\left(H_{n}(\mathbf{x}),H_{n}(\mathbf{y})\right)$ for two branches branches $\mathbf{x}=(x_{1},\dots,x_{n})$ and $\mathbf{y}=(y_{1},\dots,y_{n})$ depends exclusively on the depth at which they first diverge, denoted by $C(\mathbf{x},\mathbf{y})$:
@@ 3.1.1 Lower Bound for Time-inhomogeneous Critical Parameter
with the convention that $C(\mathbf{x},\mathbf{y})=0$ if the set is empty. For such pair of branches satisfying $C(\mathbf{x},\mathbf{y})=h$, the covariance of their Hamiltonian is $\text{Cov}\left(H_{n}(\mathbf{x}),H_{n}(\mathbf{y})\right)=2h$.
To illustrate the $L^{2}$ norm, we fix a branch $\mathbf{|x|}=n$ (noting there are expected $d^{n}$ such branches) and sum over all possible branching depths $h\in\{0,\dots,n\}$ relative to $\mathbf{x}$. For each fixed $\mathbf{x}$ and depth $h$, the number of branches $\mathbf{y}$ satisfying $C(\mathbf{x},\mathbf{y})=h$ has expected value $d^{n-h}$, corresponding to the number of branches in a sub-Galton-Watson tree with (nh)(n-h) generations.
@@
$\sum_{h=0}^{n}d^{-h}e^{-n\beta^{2}}e^{\beta^{2}(n+h)}=\sum_{h=0}^{n}e^{h(\beta^{2}-\log d)}$
@@
[Time-inhomogeneous $L^{2}$ Critical Parameter $\bar{\beta}_{2}$]
@@
where $C$ is a constant depending only on $\beta$. For any fixed $\beta>\sqrt{\log d}$, we can choose $\epsilon>0$ sufficiently samll so that $\delta>0$ becomes sufficiently small and satisfying $\log d+\beta^{2}(1-\delta)>0$, and the series diverges as $n\to\infty$. ∎
@@
Theorem 3.3 (Lower Bound for $\bar{\beta}_{c}$ and Universality Behavior).
(Phase Transition and Universality for Time-inhomogeneous BRW) Let $f:[0,1]\to[0,1]$ be a continuous, non-increasing function with $f(0)=1$. Define the normalized partition function for time-inhomogeneous weights by:
@@
We prove the $L^{1}$-convergence by dividing the branches into “good environment” and its complement, for both time-homogeneous and time-inhomogeneous environments. Using the Cameron-Martin-Girsanov formula mentioned previously, we can see that the branches in “good environment” in both cases dominates the total mass (i.e. the $L^{1}$-norm), and converges in $L^{2}$ to the same random variable.
Let $\beta<\alpha<2\beta$ be a fixed constant. For any branch $|\mathbf{x}|=n$, we define the “good environment” event for the time-homogeneous BRWs, denoted $G_{n}^{\alpha}(\mathbf{x})$ (roughly speaking, it corresponds to the environment where the Hamiltonian of a branch $\mathbf{x}$ remains “controlled”) for all $T\in\{n_{0},\ldots,n\}$ ,with $n_{0}\geq 0$ a fixed integer, by:
@@
which tends to 0 as $n_{0}\to\infty$ uniformly in $n$. A similar argument yields $\mathbb{E}[\bar{J}_{n}]=1-\bar{\epsilon}(n_{0})$ where $\bar{\epsilon}(n_{0})\to 0$ as $n_{0}\to\infty$ uniformly in $n$. While the right-hand side of (4) converges for any $\alpha>0$, we later use the good environment $G_{n}^{\alpha-\beta}(\mathbf{x})$ with $\alpha<2\beta$, justifying our restriction to $\beta<\alpha<2\beta$.
Next, we will show that when $\beta<\sqrt{2}\bar{\beta}_{2}$, $J_{n}$ and $\bar{J}_{n}$ are both uniformly bounded in $L^{2}$. Since we have seen in Theorem 3 that $\bar{\beta}_{2}=\beta_{2}$, we have that $\bar{J}_{n}$ is uniformly bounded in $L^{2}$ when $\beta<\beta_{c}$. For pairs of branches $(\mathbf{x},\mathbf{y})$ with $C(\mathbf{x},\mathbf{y})=h$, define the transformed measure $\tilde{\mathbb{P}}^{h}$ by:
@@
via Gaussian estimates. Finally, by letting $\alpha\to\beta$, we can see that the series is uniformly bounded in $n$ when $\beta<\sqrt{2}\bar{\beta_{2}}$, by a direct comparison with the $L^{2}$ computation of $\bar{W}_{n}^{\beta}$ in equation 3. A similar comparison with the $L^{2}$ computation of $W_{n}^{\beta}$ yields that when $\beta<\sqrt{2}\beta_{2}=\beta_{c}$, $J_{n}$ is uniformly bounded in $L^{2}$.
@@
Fix $n_{1}>0$, and we decompose $\mathbb{E}\left[(\bar{J}_{n}^{\beta})^{2}-\bar{J}_{n}^{\beta}J_{n}^{\beta}\right]$ by $C(\mathbf{x},\mathbf{y})=h\in\{0,...,n\}$ into two sums over $h\leq n_{1}$ and $h>n_{1}$, so that $\mathbb{E}\left[(\bar{J}_{n}^{\beta})^{2}-\bar{J}_{n}^{\beta}J_{n}^{\beta}\right]$ equals:
@@
For $\beta<\beta_{c}$ and any fixed $n_{1}>0$, form Lemma 3.4 we can see that $\bar{J}_{n}J_{n}$ and $\bar{J}_{n}^{2}$ converges to the same random variable $L^{1}$, since $\mathbb{E}[(\bar{J}_{n})^{2}]$ and $\mathbb{E}[\bar{J}_{n}J_{n}]$ has the following expansions:
@@
($L^{2}$ Convergence of Partition Function in Good Environment) Under the notation of Theorem 3.3’s proof, $J_{n}$ converges in $L^{2}$ for each fixed $n_{0}\geq 0$.
@@
(Convergence of Joint Probability in Good Environment) Let $n_{1}>0$ be a fixed integer. For all $h\in\{0,\ldots,n_{1}\}$, the convergence
@@
Remark 3.6.
Indeed, we can see from the proof that monotone decreasing and continuity is not actually relied in the proof in this section: such convergence holds as long as $f$ is continuous at 0, satisfying $f(1)=1$ and $f(x)\in[0,1]$ for all $x\in[0,1]$. This observation will be used in the proof in CREM model in Section 3.2.2, where the covariance structure could be more arbitrary.
3.1.2 Upper Bound for Time-inhomogeneous Critical Parameter
The following proposition establishes the upper bound for $\bar{\beta}_{c}$, i.e. when $\beta>\beta_{c}$, $\bar{W}_{n}^{\beta}$ converges to 0 in probability. Together with Theorem 3.3, this confirms that the normalized partition function $\bar{W}_{n}^{\beta}$ exhibits distinct phase transitions in its non-degeneracy of in-probability limit. Furthermore, the critical parameter $\bar{\beta}_{c}$ is defined by:
@@ 3.1.2 Upper Bound for Time-inhomogeneous Critical Parameter
exists and coincides with that for homogeneous weights, i.e., $\beta_{c}=\bar{\beta}_{c}$. This result underscores the insensitivity to environmental variations, indicating that the macroscopic behavior of the partition function is dominated by the “early generations” of the branching structure, rather than being affected by prescribed variations in the branch weights.
Theorem 3.7 (Upper Bound for $\bar{\beta}_{c}$).
@@
We define a transformed (random) probability measure $\bar{\mathbb{P}}^{n_{1}}$ by (where $\mathbf{P}^{*}$ is the survival probability of GW tree):
$\frac{\mathrm{d}\bar{\mathbb{P}}^{n_{1}}(\mathbf{x})}{\mathrm{d}\mathbf{P}^{*}}=\frac{e^{-\frac{1}{2}\beta^{2}\sum_{i=1}^{n_{1}}f\left(\frac{i}{n}\right)+\beta\bar{H}_{n_{1}}(\mathbf{x})}}{\bar{W}_{n}^{n_{1}}}$ (12)
@@
For any function $F:D_{n_{1}}\to\mathbb{R}$ (recall that $D_{n_{1}}$ is the set of branches $|\mathbf{x}|=n_{1}$), we have
@@
which is $\mathcal{F}_{n_{1}}$-measurable (where $\mathcal{F}_{n_{1}}$ is the $\sigma$-field generated by the GW tree and the weights up to generation $n_{1}$).
It follows that:
@@
By the Cauchy-Schwarz inequality, this implies:
@@
We will bound $\mathbb{E}\left[(\bar{W}_{n}^{n_{1}})^{2\alpha}\right]$ by:
$\mathbb{E}\left[(\bar{W}_{n}^{n_{1}})^{2\alpha}\right]\leq d^{n_{1}}\mathbb{E}\left[d^{-2\alpha n}e^{-\alpha\beta^{2}\sum_{i=1}^{n_{1}}f\left(\frac{i}{n}\right)+2\alpha\beta\bar{H}_{n_{1}}(\mathbf{x})}\right]$
@@
3.2 Applications: Time-homogeneous BRWs and Continuous Random Energy Model
In this section, we will see that the framework discussed in Section 3.1 could be naturally applied to two related models: the time-homogeneous BRWs and the Continuous random energy model (CREM).
3.2.1 Biggins Martingale Convergence: Theorem 2.1
Using the strategy of the proof of these aforementioned results, we can provide a novel proof for Theorem 2.1. Since the proof follows similar lines to that of Theorem 3.3, we prefer not to state all the details. Indeed, we can see that the proof is much simpler compared to the proof of Theorem 3.3: since $(W_{n}^{\beta})_{n\geq 0}$ defines a unit-mean martingale, to show $L^{1}$ convergence it is sufficient to show that it is uniformly integrable (i.e., it is suffice to show that $J_{n}$ is uniformly bounded in $L^{2}$ for each fixed $n_{0}>0$). The upper bound follows from the same argument as the last part of the proof of Proposition 3.7: for $\alpha\in(0,1)$ we have:
@@ 3.2 Applications: Time-homogeneous BRWs and Continuous Random Energy Model
3.2.2 Convergence of the Partition Function of CREM: Proposition 2.4
Here, we will provide an extension of our result to the continuous random energy model (CREM), see Section 2 for the precise definition of this model.
In [6], Bovier and Kurkova showed that there exists a critical parameter $\beta_{c}=\frac{\sqrt{2\log 2}}{\sqrt{\hat{A}^{\prime}(0)}}$ such that the following phase transition occurs: a) For all $\beta<\beta_{c}$, the main contribution to the partition function comes from an exponential number of the particles. b) For all $\beta>\beta_{c}$, the maximum starts to contribute significantly to the partition function. The quantity $\beta_{c}$ is sometimes referred to as the static critical inverse temperature of the CREM, and the coincidence with $\bar{\beta}_{c}$ (if we normalize $A$ so that $\hat{A}^{\prime}(0)=1$) in the present paper further suggests the potential relationship between these two models. Under the framework of the present paper, we can see that although the exact approximation terms are different in this setting, the Gaussian increment and the covariance structure allow us to prove the following:
Proposition 3.8.
(Convergence of CREM Partition Function) Suppose for any $x\in[0,1]$, $A^{\prime}(x)$ exists almost everywhere, satisfies $\sup_{x\in[0,1]}A^{\prime}(x)\leq A(0)$ and continuous at 0 (i.e., $A\in C^{0}[0,1]\cap C^{1}(0)$). Then when $\beta<\beta_{c}=\frac{\sqrt{2\log 2}}{\sqrt{A^{\prime}(0)}}$, we have $Z_{n}^{\beta}$ converges to $Z_{\infty}^{\beta}$ in $L^{1}$ and $Z_{\infty}^{\beta}$ has the same law as $W_{\infty}^{\beta}$, and when $\beta>\beta_{c}$, we have $Z_{n}^{\beta}$ converges to 0 in probability.
Proof.
(Sketch) Indeed, we can easily see that the proof follows similar lines to that of Theorem 3.3, so we only sketch the proof here. Under the binary tree at generation $n$ and its covariance structure, for each branch $|\mathbf{x}|=n$, and any $0\leq k\leq N$, we can see that the Hamiltonian of $u$ at generation $k$ has law $\mathcal{N}\left(0,NA(\frac{k}{N})\right)$. As a result, the $L^{2}$ norm of $Z_{n}^{\beta}$ is given by:
$\mathbb{E}\left[\left(Z_{n}^{\beta}\right)^{2}\right]=\sum_{h=0}^{n}2^{-h}e^{\beta^{2}nA(\frac{h}{N})}$
Since $A^{\prime}(x)$ is continuous at $x=0$, by applying a similar argument to that in Theorem 3.1 we can see that it is uniformly bounded when $\beta<\frac{\sqrt{\log 2}}{\sqrt{A^{\prime}(0)}}$. Then we may apply the same measure tilting argument and obtain the exact law of the Hamiltonian $H_{n}(\mathbf{x})$ (where for each $T\in\{0,\dots,n\}$ and $|\mathbf{x}|=T$, $H_{T}(\mathbf{x})\sim\mathcal{N}(0,A(\frac{T}{N}))$) under the tilted measure, similarly to that in Theorem 3.3, and the rest of the proof follows since the continuity of $A^{\prime}$ at 0 implies that for any finite $n_{0}>0$, we have $A(\frac{T}{N})\to A(0)$ for any $0\leq T\leq n_{0}$.
The upper bound for $\beta_{c}$ follows from the same lines as in Theorem 3.7. ∎
3.3 Behavior of Partition Function at Criticality